THE ROTATIONAL SPECTRUM OF O_{2}-HCl

\underline{TETSUYA MORI}, YOSHIHIRO SUMIYOSHI and YASUKI ENDO, Department of Basic Science, Graduate School of Arts and Sciences, The University of Tokyo, Komaba, Meguro-ku, Tokyo, 153-8902, Japan.

Rotational spectra of O_{2}-H^{35}Cl and O_{2}-H^{37}Cl have been studied from 3.5 to 21 GHz with an FTMW spectrometer in combination with a supersonic beam system. Several series of lines have been observed in the region. One series consists of strong lines with about 3800 MHz interval, starting at the lowest transition, J=1-0. Lines of this series have a small Zeeman effect in comparison with other series. All the observed lines have hyperfine structures caused by the quadrupole coupling of Cl and a much smaller magnetic hyperfine coupling of the proton. The molecular structure, results of ab initio calculations, and the rotational and hyperfine constants will be discussed.